I'm not sure if i selected the correct type of question for the subject but I need help on a "Number Problem"
The book I have does not explain how to solve, it just has them in the back as a question.
The sum of two numbers is 27 and there difference is 20. Find the numbers.
How would you start this problem?
: I'm not sure if i selected the correct type of question for the subject but I need help on a "Number Problem"
The book I have does not explain how to solve, it just has them in the back as a question.
The sum of two numbers is 27 and there difference is 20. Find the numbers.
How would you start this problem?

Answer by vleith(964)

This is a classic algebra problem.
Let the two numbers be represented as x and y.
You are given:
x + y = 27 the sum is 27
x - y = 20 their difference is 20
You can use substitution or elimination to solve these. See http://collabvsl.wetpaint.com/page/Solving+Systems+of+Equations for a nice video that covers different ways to solve systems like this.
We'll use elimination to solve
x + y = 27
x - y = 20 add these two equations
-----------------
2x + 0y = 47
2x = 47
x = 47/2
x = 23.5

Now that you know one number, use that to solve for the other one
x+y = 27
23.5 + y = 27
y = 27 - 23.5
y = 3.5
You can verify the answers by substituting back into the original two equations.